0 0
Add a comment Improve this question Transcribed image text
Answer #1

Given : \lambda= 470 nm ; d = 1.55×10-3cm = 0.0155×10-3 m

Solution:

we know that , sin\theta = m\lambda/d

sin\theta = (2)(470×10-9m)/(0.0155×10-3m)

sin\theta = 0.060645

i.e. \theta = sin-1(0.060645)= 3.48°